OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? A. 25, 23, 7 B. 9, 6, 3 C. 18, 15, 4 D. none of the above

OpenStudy (anonymous):

D

OpenStudy (welshfella):

for a right angled triangle the square on the longest side = sum of the squares on the other 2 sides

OpenStudy (anonymous):

A right angle equals 90 degrees... add them the numbers up to see which answer has the sum of 90 degrees

OpenStudy (welshfella):

if its D then 18^2 = 15^2 + 4^2 check that out

OpenStudy (anonymous):

i did @welshfella

OpenStudy (welshfella):

and what did you get?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

for the final answer because i had 324 and took the square root of that and got 18

OpenStudy (welshfella):

?? did 18^2 = 15^2 + 4^2?

OpenStudy (anonymous):

its none of the above try them all none will work

OpenStudy (welshfella):

18^2 = 324 15^2 = 225 4^2 = 16 now does the sum of the last 2 equal 324?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

?? 225 + 16 = 241 - not 324

OpenStudy (anonymous):

yes thats what i got not 324 because i wrote it out on scratch paper and had that the second time. the first time i had gotten 324

OpenStudy (welshfella):

do you see that this cannot be a right angled triangle because the numbers dont fit the pythagoras theorem?

OpenStudy (anonymous):

yes, so it would be D

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Thank You! :)

OpenStudy (anonymous):

Wanna help me with another one?

OpenStudy (welshfella):

sorry - I made a mistake when I said if its D - i should have said if its C (in my second post)

OpenStudy (welshfella):

yea ok

OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? 101, 99, 20 28, 26, 12 17, 14, 6 none of the above

OpenStudy (anonymous):

Add them

OpenStudy (anonymous):

For example : Add 100+99+20 and see if it gets you 90

OpenStudy (welshfella):

again you have to go throgh each one and see if they fit pythagoras eg 100^2 = 99^2 + 20^2 ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so if i get 90 as my final answer then it is the answer?

OpenStudy (welshfella):

no - its not adding the numbers - you add the SQUARES of the numbers

OpenStudy (anonymous):

yes. Add the squares

OpenStudy (anonymous):

how do i know which is the hypotenuse?

OpenStudy (welshfella):

its the longest side

OpenStudy (welshfella):



OpenStudy (welshfella):

i'll do one for you 100^2 = 10000 99^2 = 9801 20^2 = 400 9801 + 400 = 10201 so its not a right angle triangle

OpenStudy (anonymous):

i did the first one i added the square of 99 and got 9801 then added that with the square of 20 which was 400 and together those equaled 10201 and the sum of 101 squared was 10201

OpenStudy (welshfella):

yes good

OpenStudy (anonymous):

okay so the next one now

OpenStudy (welshfella):

so the first one cannot be right angled

OpenStudy (welshfella):

now do the same calculations on the other 2

OpenStudy (anonymous):

okay brb

OpenStudy (anonymous):

so i got 820 for the sqaure of 26 and square of 12. Then i got 784 for the square of 28

OpenStudy (welshfella):

`right

OpenStudy (welshfella):

so is the second set a right angled triangle or not?